OpenStudy (anonymous):

Find the absolute extrema of the function on the closed interval. f(x) = x^3 − (3/2)x^2 [−1, 6] minimum (x,y) = maximum (x,y) =

OpenStudy (mathmale):

This problem has two parts. One part is to determine any relative minima/maxima within the interval [-1,6]. The other is to determine the values of the given function at the endpoints. Which of these are you able to do now?

OpenStudy (anonymous):

I differentiated the function as 3x^2-2(3/2)x This became 3x(1-x)

OpenStudy (mathmale):

Agreed, except that I'd reverse the quantity within parentheses: 3*x*(x-1). Can you agree with that?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Ok. that's your derivative. set this = to 0 and solve for y our critical values. What are they?

OpenStudy (anonymous):

3x(x-1)=0 x=0,1 f(0)= -1 f(1)=0 f(-1)=6 f(6)=90 critical values = -1,90

OpenStudy (mathmale):

Your critical values are actually {0,1}. These are the roots of the equation f '(x)=0, right? When i evaluated f(x) = x^3 − (3/2)x^2 at -1, I got -2.5; at 6, I got 162. Would you mind double-checking your work there?

OpenStudy (mathmale):

Is this correct?\[f(x) = x^3 − (3/2)x^2 \]

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Then\[f(x) = (-1)^3 − (3/2)(-1)^2 =?\]

OpenStudy (anonymous):

2.25

OpenStudy (mathmale):

Here's what I get: f(-1)=-1-(3/2)(1)=-2.5.

OpenStudy (mathmale):

We need to agree on all the y-values before we can proceed.

OpenStudy (mathmale):

\[f(6) = (6)^3 − (3/2)(6)^2 =?\]

OpenStudy (anonymous):

162

OpenStudy (mathmale):

I agree with that. Let's now make a table: